OpenStudy (anonymous):

A cotton candy holder is shaped like a cone. The height is 13 in., and the diameter is 8 in. What is the volume of the holder? Use 3.14 to approximate pi, and round answers to two decimal places if necessary. _____ in3

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

1/3 pi x r^2 x h

OpenStudy (anonymous):

am i doing this right 8 times 8 is 64 and 64 times 13 is 832

OpenStudy (anonymous):

and 832 divide by 3 is 277

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@kliedako

OpenStudy (sleepyjess):

Pretty close, we are given the diameter, which is r*2. So to find the radius, we need to do d/2

OpenStudy (anonymous):

8/2?

OpenStudy (anonymous):

@BrownieWuvers @sleepyjess got this

OpenStudy (anonymous):

okay @kliedako

OpenStudy (sleepyjess):

Yes, so \(\dfrac{4^2*h*\pi}3\)

OpenStudy (sleepyjess):

and the height, h, is given as 13

OpenStudy (anonymous):

69.3?

OpenStudy (sleepyjess):

How did you get to that?

OpenStudy (anonymous):

4 times 4 = 8 and 8 times 13 is 104

OpenStudy (anonymous):

oh. then 34.6

OpenStudy (sleepyjess):

4*4 = 8?

OpenStudy (anonymous):

-.- my bad

OpenStudy (anonymous):

16

OpenStudy (anonymous):

i'm really tired and i was adding lol .-.

OpenStudy (sleepyjess):

haha :)

OpenStudy (sleepyjess):

Ok, so now we have (16*13*3.14)/3

OpenStudy (anonymous):

653 :)

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

yes, now 653/3

OpenStudy (anonymous):

217

OpenStudy (sleepyjess):

yep

OpenStudy (anonymous):

Thank you sleepy jes cx

OpenStudy (sleepyjess):

no problem! :)

OpenStudy (anonymous):

Have a nice night lol

OpenStudy (anonymous):

The answer was 217.71.

OpenStudy (anonymous):

:"(

OpenStudy (sleepyjess):

:(